# Random-sketching Techniques to Enhance the Numerically Stability of Block Orthogonalization Algorithms for $s$-step GMRES

### v1

@@ Abstract
We integrate random sketching techniques into block orthogonalization schemes needed for $s$-step GMRES. The resulting block orthogonalization schemes generate the basis vectors whose overall orthogonality error is bounded by machine precision as long as each of the corresponding block vectors are numerically full rank. We implement these randomized block orthogonalization schemes using standard distributed-memory linear algebra kernels for $s$-step GMRES available in the Trilinos software packages. Our performance results on the Perlmutter supercomputer (with four NVIDIA A100 GPUs per node) demonstrate that these randomized techniques can enhance the numerical stability of the orthogonalization and overall solver, without a significant increase in the execution time.
@@ 1 Introduction
Generalized Minimum Residual (GMRES) [1] is a popular subspace projection method for iteratively solving a large linear system of equations as it computes the approximate solution that minimizes the residual norm in the generated Krylov subspace. To compute the approxmate solution, GMRES generates the orthogonal basis vectors of the projection subspace based on two main computational kernels: 1) Sparse-Matrix Vector multiply (SpMV), typically combined with a preconditioner, and 2) orthogonalization. Though GMRES is a robust iterative method for solving general linear systems, the performance of these two kernels can be limited by communication costs (e.g., the cost of moving data through the local memory hierarchy and between the MPI processes). For instance, on a distributed-memory computer, to orthogonalize a new basis vector at each iteration, GMRES requires global reduces among all the MPI processes and performs its local computation based on either BLAS-1 or BLAS-2 operations. Hence, though the breakdown of the iteration time depends on the preconditioner being used, orthogonalization can become a significant part of the iteration time, especially when scalable implementations of SpMV and preconditioner are available.
@@
Since the potential speedup from the block orthogonalization is limited by the small step size $s$ required to maintain the numerical stability of MPK, a two-stage variant of block orthogonalization was proposed [6]. In order to maintain the well-conditioning of the basis vectors, the first stage of this two-stage algorithm pre-processes the block of $s+1$ basis, while the full orthogonalization is delayed until enough number of basis vectors, $\widehat{s}+1$, are generated to obtain the higher performance. This improves the performance of the block orthogonalization process while using the small step size $s$.
Block orthogonalization consists of inter and intra block orthogonalization to orthogonalize a new block of vectors against the already-orthogonalized blocks of vectors and among the vectors within the new block, respectively. For the inter-block orthogonalization, Block Classical Gram-Schmidt with re-orthogonalization (BCGS twice, or BCGS2) obtains good performance on current hardware architectures because it is based entirely on BLAS-3. For robustness and performance of the overall block orthogonalization, the critical component is the algorithm used for the first intra-block orthogonalization [7]. In this paper, we consider the use of CholQR [8] twice (CholQR2) as our intra-block orthogonalization, which is based mainly on BLAS-3. Unfortunately, though the above combinations of the algorithms performs well on current hardware, $s$-step basis vectors can be ill-conditioned, and CholQR2 can fail when the condition number of $s+1$ vectors is greater than the reciprocal of the square-root of machine epsilon.
@@
    We integrate random sketching techniques into BCGS2 such that overall orthogonalization error is on the order of machine precision in both one-stage and two-stage frameworks as long as the corresponding block vectors are numerically full-rank.
@@
    We present numerical results to demonstrate the improved numerical stability using random sketching techniques (to pre-process the basis vectors) compared to the state-of-the-art deterministic algorithms (BCGS2 with CholQR).
@@
Table 1 lists the notation used in this paper. In addition, we use $Q_{\ell:t}$ to denote the blocks column vectors of $Q$ with the block column indexes $\ell$ to $t$, while $q_{k:s}$ is the set of vectors with the column indexes $k$ to $s$. We then use the bold small letter $\mathbf{v}_{j}$ to denote the sketched version of the block vector $V_{j}$, e.g., $\mathbf{v}_{j}=\Theta^{T}V_{j}$. Finally, $[Q,V]$ is the column concatenation of $Q$ and $V$. In addition, $c_{k}(\epsilon,n,s)$ is a scalar function that depends on $n$, $\epsilon$, and $s$.
@@ 2 Related Work
    One reason for this is that except for one special case (i.e., two-stage with $\widehat{s}=m$, where $m$ is the restart length), we sketch only a part of the Krylov subspace (e.g., each panel or big panel of $s+1$ or $\widehat{s}+1$ basis vectors, respectively), requiring the $\ell_{2}$-orthogonality for the consistency over the restart loop.
@@
    In particular, we focus on enhancing the numerical stability of the solver using such random sketching techniques. Specifically, we develop randomized block orthogonalization algorithms that obtain $O(\epsilon)$ orthogonality errors, given the corresponding block vectors are numerically full-rank.
  • 3.
@@
We focus on the practical implementations of random sketching, using standard linear algebra kernels, in particular its effects on the performance of the $s$-step GMRES on a GPU cluster. The GPU performance of random sketching [19, 20] has been previously studied. This paper differs from these previous works since we focus on the algorithmic development of numerically stable block orthogonalization schemes. We then integrate the random sketching techniques into $s$-step GMRES, utilizing the readily-available computational kernels, and study their impact on the distributed GPU performance of $s$-step GMRES.
@@ 3 Background
By $\Theta\in\mathbb{R}^{n\times\widehat{m}}$, we denote a random sketch matrix, with $\widehat{m}\ll n$. The general concept of “random sketching” is to apply this random sketch matrix $\Theta$ (generated using some randomization strategy) to a large matrix $V$ to obtain a “sketched” matrix $\mathbf{v}=\Theta^{T}V$ that greatly reduces the row dimension of $V$ while preserving its fundamental properties, such as its norm and singular values, as much as possible. In particular, the sketch matrix $\Theta$ is typically chosen to be a subspace embedding, or a linear map to a lower dimensional space that preserves the $\ell_{2}$-inner product of all vectors within the subspace up to a factor of $\sqrt{1\pm\varepsilon}$ for some $\varepsilon\in(0,1)$ [21, 22]. Such embeddings also preserve $\ell_{2}$-norms in a similar way [9].
Definition 3.1 ($\varepsilon$-subspace embedding).
Given $\varepsilon\in(0,1)$, the sketch matrix $\Theta\in\mathbb{R}^{n\times\widehat{m}}$ is an $\varepsilon$-subspace embedding for the subspace $\mathcal{V}\subset\mathbb{R}^{n}$ if $\forall x,y\in\mathcal{V}$,
$|\langle x,y\rangle-\langle\Theta^{T}x,\Theta^{T}y\rangle|\leq\varepsilon\|x\|_{2}\|y\|_{2}.$ (1)
@@
If the sketch matrix $\Theta\in\mathbb{R}^{n\times\widehat{m}}$ is an $\varepsilon$-subspace embedding for the subspace $\mathcal{V}\subset\mathbb{R}^{n}$, then $\forall x\in\mathcal{V}$,
$\sqrt{1-\varepsilon}~{}\|x\|_{2}\leq\|\Theta^{T}x\|_{2}\leq\sqrt{1+\varepsilon}~{}\|x\|_{2}.$ (2)
@@
If the sketch matrix $\Theta\in\mathbb{R}^{n\times\widehat{m}}$ is an $\varepsilon$-subspace embedding for the subspace $\mathcal{V}\subset\mathbb{R}^{n}$, and $V$ is a matrix whose columns form a basis of $\mathcal{V}$, then
$(1+\varepsilon)^{-1/2}~{}\sigma_{min}(\Theta^{T}V)$ $\leq\sigma_{min}(V)\leq\sigma_{max}(V)$ (3)
$\leq(1-\varepsilon)^{-1/2}~{}\sigma_{max}(\Theta^{T}V).$
@@
$\kappa(V)\leq\sqrt{\frac{1-\varepsilon}{1+\varepsilon}}~{}\kappa(\Theta^{T}V).$ (4)
@@
The limitation of $\varepsilon$-subspace embedding presented in Definition 3.1 is that to ensure that the sketch matrix approximately preserves norms and inner products, one needs to know the subspace $\mathcal{V}\subset\mathbb{R}^{n}$ a priori. In contrast, to use sketching techniques in Krylov subspace methods efficiently, we need a sketch matrix that does not require complete prior knowledge of the subspace, since Krylov subspaces are generated as the algorithm iterates. This can be accomplished by using $(\varepsilon,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embeddings [9].
Definition 3.2 ($(\varepsilon,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embedding).
The sketch matrix $\Theta\in\mathbb{R}^{n\times\widehat{m}}$ is an $(\varepsilon,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embedding if it is an $\varepsilon$-subspace embedding for any fixed $\widehat{s}$-dimensional subspace $\mathcal{V}\subset\mathbb{R}^{n}$ with probability at least $1-\delta$.
Given a Gaussian matrix $G\in\mathbb{R}^{n\times\widehat{m}}$, a concrete example of a $(\varepsilon,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embedding is $\Theta=\frac{1}{\sqrt{\widehat{m}}}G$ where $\widehat{m}=\Omega(\varepsilon^{-2}\log\widehat{s}\log(1/\delta))$ [22]. A far simpler relation $\widehat{m}\approx\widehat{s}/\varepsilon^{2}$ has been also proposed for an “optimal scaling” of the embedding length [21]. This relation allows a simple correspondence between the embedding dimension (sketch size) $\widehat{m}$ and the subspace dimension $\widehat{s}$ for a given $\varepsilon$. For instance, to achieve $\varepsilon=1/\sqrt{2}$, the sketch size of $\widehat{m}\approx 2\widehat{s}$ is sufficient, though in principle, one could choose a different value of $\varepsilon$ to construct a different sketch size. Other $(\varepsilon,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embeddings exist that can be stored in a sparse format, including sub-sampled randomized Hadamard and Fourier transforms (SRHT and SRFT, respectively), and “sparse dimension reduction maps” [9, 21].
@@ 4 Block Orthogonalization for $s$-step GMRES
Compared to the standard GMRES, $s$-step GMRES has the potential to reduce the communication cost of generating the $s$ basis vectors by a factor of $s$, where standard GMRES is essentially $s$-step GMRES with the step size of one. For instance, to apply SpMV $s$ times (Lines 7 to 9 of the pseudocode), several CA variants of MPK exist [25]. On a distributed-memory computer, CA variants may reduce the communication latency cost of SpMV, associated with the point-to-point neighborhood Halo exchange of the input vector, by a factor of $s$ (though it requires additional memory and local computation, and it may also increase the total communication volume). However, in practice, SpMV is typically combined with a preconditioner to accelerate the convergence rate of GMRES, but only a few CA preconditioners of specific types have been proposed [26, 27]. To support a wide-range of preconditioners used by applications, instead of CA MPK, Trilinos $s$-step GMRES uses a standard MPK (applying each SpMV with neighborhood communication in sequence), and focuses on improving the performance of block orthogonalization. Also, avoiding the global communication may lead to a greater gain on the orthogonalization performance than CA MPK does on SpMV performance.
@@
$\|I-\widehat{Q}_{j}^{T}\widehat{Q}_{j}\|$ $=$ $\mathcal{O}(\epsilon).$ (6)
In the next section, we explore two algorithms for the first intra-block orthogonalization. Our focus is on algorithms that perform well on current hardware architectures and on the condition on each block vector (i.e., the required condition number of the basis vectors, $\kappa(\widehat{V}_{j})$) that is sufficient in order for each algorithm to guarantee (6).
@@ 5.1 CholQR twice (CholQR2)
Overall, the orthogonality error of $Q_{j}$ generated by CholQR is bounded as
@@
and BCGS2 with CholQR2 obtains the $\mathcal{O}(\epsilon)$ orthogonality error when the following condition is satisfied [6, 8, 30, Theorem IV.1].
Proposition 5.1.
@@
the orthgonality error of the basis vectors $\widehat{Q}_{j}$ computed by CholQR2, and hence of the basis vectors $Q_{1:j}$ generated by BCGS2 with CholQR2, is of the order of the machine precision (as the condition (6) ensures the condition (5)).
The proof of Proposition 5.1 follows by combining the results from two facts: first, BCGS2 attains $O(\epsilon)$ orthogonality error overall provided each “IntraBlk” step in line 3 of Figure 2(c) attains $O(\epsilon)$ orthogonality error [7], and CholQR2 attains $O(\epsilon)$ orthogonality error when (8) is satisfied [31].
The main drawback of CholQR is that it computes the Gram matrix of the input basis vectors to be orthogonalized (Line 2 in Figure 2(b)), and the Gram matrix has the condition number which is the square of the input vectors’ condition number. Hence, CholQR can fail when the condition number of the input vectors $V_{j}$ is greater than the reciprocal of the square-root of the machine epsilon $\epsilon$ (i.e., $\kappa(V_{j})>1/\mathcal{O}(\epsilon^{1/2})$). This can cause numerical issues, especially for the $s$-step method, because even when Newton or Chebyshev basis are generated, the $s$-step basis vectors can be ill-conditioned with a large condition number.
To alleviate this potential numerical instability, Trilinos implements a “recursive” variant of CholQR, as shown in Figure 3; when Cholesky factorization of the Gram matrix fails at the $k$th step due to a non-positive diagonal, it orthogonalizes just the first $k-1$ vectors by CholQR. It then orthogonalizes the remaining vectors against the first $k-1$ orthonormal vectors by BCGS and recursively calls CholQR on the remaining vectors. This avoids the algorithmic breakdown of the orthogonalization process by adaptively adjusting the block size to orthogonalize the $s+1$ basis vectors. To orthogonalize the remaining vectors against the first $k-1$ vectors, it uses the partial Cholesky factors, and hence no additional overhead is needed. However, it needs to re-compute the dot-products of the remaining vectors, and may require multiple global reduces for ill-conditioned basis vectors. Furthermore, though it is often effective in recovering from the failures in combination with MPK, there is no bound on the orthogonality error with this recursive variant.

### v2

# Random Sketching to Enhance the Numerical Stability of Block Orthogonalization Algorithms for $s$-step GMRES

@@ Abstract
We integrate random sketching techniques into block orthogonalization schemes needed for $s$-step GMRES. The resulting, one-stage and two-stage, block orthogonalization schemes generate the basis vectors whose overall orthogonality error is bounded by machine precision as long as each of the corresponding block vectors are numerically full rank. We implement these randomized block orthogonalization schemes using standard distributed-memory linear algebra kernels for $s$-step GMRES available in the Trilinos software packages. Our performance results on the Perlmutter supercomputer (with four NVIDIA A100 GPUs per node) demonstrate that these randomized techniques can enhance the numerical stability of the orthogonalization and overall solver, without a significant increase in the execution time.
@@ 1 Introduction
Generalized Minimum Residual (GMRES) [1] is a popular subspace projection method for iteratively solving a large linear system of equations as it computes the approximate solution that minimizes the residual norm in the generated Krylov subspace. To compute the approxmate solution, GMRES generates the orthonormal basis vectors of its projection subspace based on two main computational kernels: 1) Sparse-Matrix Vector multiply (SpMV), typically combined with a preconditioner, and 2) orthogonalization.
Though GMRES is a robust iterative method for solving general linear systems, the performance of these two kernels can be limited by communication costs (e.g., the cost of moving data through the local memory hierarchy and between the MPI processes). For instance, on a distributed-memory computer, to orthogonalize a new basis vector at each iteration, GMRES requires global reduces among all the MPI processes and performs its local computation based on either BLAS-1 or BLAS-2 operations. Hence, though the breakdown of the iteration time depends on the target hardware architecture and problem properties (e.g., the sparsity structure of the matrix and the preconditioner being used), orthogonalization can become a significant part of the iteration time, especially when scalable implementations of SpMV and preconditioner are available.
@@
Since the potential speedup from the block orthogonalization is limited by the small step size $s$ required to maintain the numerical stability of MPK, a two-stage variant of block orthogonalization was proposed [6]. In order to maintain the well-conditioning of the basis vectors, the first stage of this orthogonalization scheme pre-processes the block of $s+1$ basis, while the full orthogonalization is delayed until enough number of basis vectors, $\widehat{s}+1$, are generated to obtain the higher performance. This improves the performance of the block orthogonalization process while using the small step size $s$.
Block orthogonalization consists of 1) inter block orthogonalization to orthogonalize a new block of vectors against the already-orthogonalized blocks of vectors and 2) intra block orthogonalization to orthogonalize among the vectors within the new block. For the inter-block orthogonalization, Block Classical Gram-Schmidt with re-orthogonalization (BCGS twice, or BCGS2) obtains good performance on current hardware architectures because it is based entirely on BLAS-3. For robustness and performance of the overall block orthogonalization, the critical component is the algorithm used for the first intra-block orthogonalization [7]. In this paper, we consider the use of CholQR [8] twice (CholQR2) as our intra-block orthogonalization, which is based mainly on BLAS-3. Unfortunately, though the above combinations of the algorithms performs well on current hardware, $s$-step basis vectors can be ill-conditioned, and CholQR2 can fail when the condition number of $s+1$ vectors is greater than the reciprocal of the square-root of machine epsilon.
@@
    We integrate random sketching techniques into BCGS2 such that overall orthogonalization error is on the order of machine precision in both one-stage and two-stage frameworks as long as each of the corresponding block of $s+1$ and $\widehat{s}+1$ vectors are numerically full-rank, respectively.
@@
    We present numerical results to demonstrate the improved numerical stability using random sketching techniques (to pre-process the basis vectors) compared to the state-of-the-art deterministic algorithms (BCGS2 with CholQR2).
@@
Table 1 lists the notation used in this paper. In addition, we use $Q_{\ell:t}$ to denote the blocks column vectors of $Q$ with the block column indexes $\ell$ to $t$, while $q_{k:s}$ is the set of vectors with the column indexes $k$ to $s$. We then use the bold small letter $\mathbf{v}_{j}$ to denote the sketched version of the block vector $V_{j}$, e.g., $\mathbf{v}_{j}=\Theta^{T}V_{j}$. Finally, $[Q,V]$ is the column concatenation of $Q$ and $V$. For our numerical analysis, we use $c_{k}(\epsilon,n,s)$ to represent a scalar constant that is in the order of the machine epsilon $\epsilon$ but also depends on the matrix dimensions $n$ and $s$.
@@ 2 Related Work and Our Motivations
    One reason for this is that except for one special case (i.e., two-stage with $\widehat{s}=m$, where $m$ is the restart length), we sketch only a part of the Krylov subspace (e.g., each panel or big panel of $s+1$ or $\widehat{s}+1$ basis vectors, respectively), requiring the $\ell_{2}$-orthogonality to ensure the overall consistency of the basis vectors over the restart loop.
@@
The GPU performance of random sketching has been previously studied [19, 20]. This paper differs from these previous works since we focus on the algorithmic development of numerically stable block orthogonalization schemes. We then integrate the resulting randomized block orthogonalization schemes into $s$-step GMRES, and study their performance impact on a GPU cluster. We focus on the practical implementations of random sketching, using readily-available standard linear algebra kernels, though specialized kernels may improve the performance.
Some of the sparse sketching techniques have the potential to reduce the computational complexity of the orthogonalization process. Unfortunately, we have not seen this performance benefit in our performance experiments, using the standard linear algebra kernels. Nervelessness, our main focus is to enhance the numerical stability of the solver by integrating the random sketching techniques. Specifically, we develop randomized block orthogonalization algorithms that obtain $O(\epsilon)$ orthogonality errors, given the corresponding block vectors are numerically full-rank.
@@ 3 Background
By $\Theta\in\mathbb{R}^{n\times\widehat{m}}$, we denote a random sketch matrix, with $\widehat{m}\ll n$. The general concept of “random sketching” is to apply this random sketch matrix $\Theta$ (generated using specific probability distributions) to a large matrix $V$ to obtain a “sketched” matrix $\mathbf{v}=\Theta^{T}V$ that greatly reduces the row dimension of $V$ while preserving its fundamental properties, such as its norm and singular values, as much as possible. In particular, the sketch matrix $\Theta$ is typically chosen to be a subspace embedding, or a linear map to a lower dimensional space that preserves the $\ell_{2}$-inner product of all vectors within the subspace up to a factor of $\sqrt{1\pm\mu}$ for some $\mu\in(0,1)$ [21, 22]. Such embeddings also preserve $\ell_{2}$-norms in a similar way [9].
Definition 3.1 ($\mu$-subspace embedding).
Given $\mu\in(0,1)$, the sketch matrix $\Theta\in\mathbb{R}^{n\times\widehat{m}}$ is a $\mu$-subspace embedding for the subspace $\mathcal{V}\subset\mathbb{R}^{n}$ if $\forall x,y\in\mathcal{V}$,
$|\langle x,y\rangle-\langle\Theta^{T}x,\Theta^{T}y\rangle|\leq\mu\|x\|_{2}\|y\|_{2}.$ (1)
@@
If the sketch matrix $\Theta\in\mathbb{R}^{n\times\widehat{m}}$ is a $\mu$-subspace embedding for the subspace $\mathcal{V}\subset\mathbb{R}^{n}$, then $\forall x\in\mathcal{V}$,
$\sqrt{1-\mu}~{}\|x\|_{2}\leq\|\Theta^{T}x\|_{2}\leq\sqrt{1+\mu}~{}\|x\|_{2}.$ (2)
@@
If the sketch matrix $\Theta\in\mathbb{R}^{n\times\widehat{m}}$ is a $\mu$-subspace embedding for the subspace $\mathcal{V}\subset\mathbb{R}^{n}$, and $V$ is a matrix whose columns form a basis of $\mathcal{V}$, then
$(1+\mu)^{-1/2}~{}\sigma_{min}(\Theta^{T}V)$ $\leq\sigma_{min}(V)\leq\sigma_{max}(V)$ (3)
$\leq(1-\mu)^{-1/2}~{}\sigma_{max}(\Theta^{T}V).$
@@
$\kappa(V)\leq\sqrt{\frac{1-\mu}{1+\mu}}~{}\kappa(\Theta^{T}V).$ (4)
@@
The limitation of $\mu$-subspace embedding presented in Definition 3.1 is that to ensure that the sketch matrix approximately preserves norms and inner products, one needs to know the subspace $\mathcal{V}\subset\mathbb{R}^{n}$ a priori. In contrast, to use sketching techniques in Krylov subspace methods efficiently, we need a sketch matrix that does not require complete prior knowledge of the subspace, since Krylov subspaces are generated as the algorithm iterates. This can be accomplished by using $(\mu,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embeddings [9].
Definition 3.2 ($(\mu,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embedding).
The sketch matrix $\Theta\in\mathbb{R}^{n\times\widehat{m}}$ is a $(\mu,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embedding if it is a $\mu$-subspace embedding for any fixed $\widehat{s}$-dimensional subspace $\mathcal{V}\subset\mathbb{R}^{n}$ with probability at least $1-\delta$.
One concrete example of a $(\mu,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embedding is $\Theta=\frac{1}{\sqrt{\widehat{m}}}G$ where $G\in\mathbb{R}^{n\times\widehat{m}}$ is a Gaussian matrix and the sketch size is given by $\widehat{m}=\Omega(\mu^{-2}\widehat{s})$ [23]. In practice, the sketch size may be chosen as $\widehat{m}\approx\widehat{s}/\mu^{2}$  [21]. This relation allows a simple correspondence between the sketch size (or equivalenty the embedding dimension) $\widehat{m}$ and the subspace dimension $\widehat{s}$ for a given $\mu$. For instance, to achieve $\mu=1/\sqrt{2}$, the sketch size of $\widehat{m}\approx 2\widehat{s}$ is sufficient, though in principle, one could choose a different value of $\mu$ to construct a different sketch size. Other $(\mu,\delta,\widehat{s})$ oblivious $\ell_{2}$-subspace embeddings exist that can be stored in a sparse format, including sub-sampled randomized Hadamard and Fourier transforms (SRHT and SRFT, respectively), and “sparse dimension reduction maps” [9, 21].
@@ 4 Block Orthogonalization for $s$-step GMRES
Compared to the standard GMRES, $s$-step GMRES has the potential to reduce the communication cost of generating the $s$ basis vectors by a factor of $s$, where standard GMRES is essentially $s$-step GMRES with the step size of one. For instance, to apply SpMV $s$ times (Lines 7 to 9 of the pseudocode), several CA variants of MPK exist [26]. On a distributed-memory computer, CA variants may reduce the communication latency cost of SpMV, associated with the point-to-point neighborhood Halo exchange of the input vector, by a factor of $s$ (though it requires additional memory and local computation, and it may also increase the total communication volume). However, while in practice, SpMV is typically combined with a preconditioner to accelerate the convergence rate of GMRES, only a few CA preconditioners of specific types have been proposed [27, 28]. To support a wide-range of preconditioners used by applications, instead of CA MPK, Trilinos $s$-step GMRES uses a standard MPK (applying each SpMV with neighborhood communication in sequence), and focuses on improving the performance of block orthogonalization. Also, avoiding the global communication may lead to a greater gain on the orthogonalization performance than CA MPK does on SpMV performance.
@@
$\|I-\widehat{Q}_{j}^{T}\widehat{Q}_{j}\|$ $=$ $\mathcal{O}(\epsilon).$ (6)
In the next section, we explore two algorithms, which performs well on the current hardware architectures and achieves (6), for the first intra-block orthogonalization. We discuss the condition on each block vector (i.e., the required condition number $\kappa(\widehat{V}_{j})$ of the input basis vectors $\widehat{V}_{j}$) that is sufficient in order for each algorithm to guarantee (6).
@@ 5.1 CholQR twice (CholQR2)
In [6, 8, 31, Theorem IV.1], it has been shown that the orthogonality error of $Q_{j}$ generated by CholQR is bounded as
@@
where $c_{1}(\epsilon,n,s)$ is a constant. Hence, BCGS2 with CholQR2 obtains the overall $\mathcal{O}(\epsilon)$ orthogonality error as formalized in the following proposition.
Proposition 5.1.
When the condition (8) is satisfied, the orthgonality error of the basis vectors $\widehat{Q}_{j}$ computed by CholQR2, and hence of the basis vectors $Q_{1:j}$ generated by BCGS2 with CholQR2, is of the order of the machine precision (as the condition (6) ensures the condition (5)).
Proof.
The proof of Proposition 5.1 follows by combining the results from two facts: first, BCGS2 attains $O(\epsilon)$ orthogonality error overall provided each “IntraBlk” step on Line 3 of Figure 2(c) attains $O(\epsilon)$ orthogonality error [7], and CholQR2 attains $O(\epsilon)$ orthogonality error when (8) is satisfied [32]. ∎
The main drawback of CholQR is that it computes the Gram matrix of the input basis vectors $V_{j}$ to be orthogonalized (Line 2 in Figure 2(b)), and the Gram matrix has the condition number which is the square of the input vectors’ condition number. Hence, CholQR can fail when the condition number of the input vectors $V_{j}$ is greater than the reciprocal of the square-root of the machine epsilon $\epsilon$ (i.e., $\kappa(V_{j})>1/\mathcal{O}(\epsilon^{1/2})$). This can cause numerical issues, especially for the $s$-step method, because even when Newton or Chebyshev basis are generated, the $s$-step basis vectors can be ill-conditioned with a large condition number.
@@
To alleviate this potential numerical instability, Trilinos implements a “recursive” variant of CholQR, as shown in Figure 3; when Cholesky factorization of the Gram matrix fails at the $k$th step due to a non-positive diagonal, it orthogonalizes just the first $k-1$ vectors by CholQR. It then orthogonalizes the remaining vectors against the first $k-1$ (roughly) orthonormal vectors by BCGS and recursively calls CholQR on the remaining vectors. This avoids the algorithmic breakdown of the orthogonalization process by adaptively adjusting the block size to orthogonalize the $s+1$ basis vectors. To orthogonalize the remaining vectors against the first $k-1$ vectors, it uses the partial Cholesky factors, and hence no additional overhead is needed. However, it needs to re-compute the dot-products of the remaining vectors, and may require multiple global reduces for ill-conditioned basis vectors. Furthermore, though it is often effective in recovering from the failures in combination with MPK, there is no bound on the orthogonality error with this recursive variant.